OpenStudy (anonymous):

Does anyone know how to add and subtract rational expressions? help very last assignment D:

OpenStudy (anonymous):

Its an assignment I have to do, but its all about adding and subtracting rational expression, only 5 problems

OpenStudy (nurali):

Hope this will help

OpenStudy (anonymous):

Ok c: Thanks

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

is it ok if i do one problem and show you to make sure im not doing anything wrong?

OpenStudy (anonymous):

@Nurali

OpenStudy (usukidoll):

post an example

OpenStudy (usukidoll):

I think the other user is afk or something

OpenStudy (anonymous):

ill draw a picture of it

OpenStudy (anonymous):



OpenStudy (usukidoll):

hmmm... they share a common denominator so all we have to do is to combine like terms in the numerator

OpenStudy (anonymous):

How would we do that? would we add 2x to x and 1 to the other 1?

OpenStudy (usukidoll):

yeah

OpenStudy (anonymous):

would the numerator come out like this = 2x^2+2 or 3x+2

OpenStudy (usukidoll):

it would be 3x+2. because 2x+x = 3x 1 + 1 =2

OpenStudy (anonymous):

Ok! thanks, can you help with four more adding/subracting q's?

OpenStudy (usukidoll):

sure

OpenStudy (anonymous):

Thankyou so much! so the next question basically wants me to use the answer we just got (3x+2/2x) subtracted by x/x+3, but it also says to * identify any possible restrictions that exist with (or in) the resulting rational expression*

OpenStudy (usukidoll):

\[\frac{3x+2}{2x}-\frac{x}{x+3}\] it said any possible restrictions... so I"m not sure what that means 100% maybe what value makes the result undefined ?>! Not sure.. any way our lcd = (2x)(x+3) so we multiply x+3 on the first fraction and multiply 2x on the second fraction

OpenStudy (anonymous):

so like cross multiplying?

OpenStudy (usukidoll):

umm... similar but instead we need to find the least common denominator see example 3 on the cliffnotes site... noticed how they took the lcd of that problem?

OpenStudy (anonymous):

It looks kinda confusing sine the problem big :x

OpenStudy (anonymous):

on cliff notes

OpenStudy (usukidoll):

\[\frac{3x+2}{2x} \times \frac{x+3}{x+3}-\frac{x}{x+3} \times \frac{2x}{2x}\]

OpenStudy (anonymous):

ok that makes alot more since lol

OpenStudy (usukidoll):

example 3 on cliffnotes from Nurali was just x and y as the denominators. it was the later examples that I didn't choose because I knew it would be too much

OpenStudy (anonymous):

would we just times the bottom demonitaers togethor? like 2x(x+3)

OpenStudy (usukidoll):

yes...

OpenStudy (usukidoll):

\[\frac{(3x+2)(x+3)-2x^2}{2x(x+3)}\] we have to remember that whenever we have a - sign on the right side of the equation, we need to distribute that sign all over

OpenStudy (usukidoll):

\[-x(2x) \rightarrow -2x^2\]

OpenStudy (usukidoll):

and we need to use FOIL method to expand the left side

OpenStudy (anonymous):

I know how to do that cx ill show you after im done foiling it c:

OpenStudy (usukidoll):

ok

OpenStudy (anonymous):

would we do the distributive property to the denomater? or does that stay the same

OpenStudy (usukidoll):

umm I think it depends on the numerator... because we wouldn't want to combine the denominator and end up that we can cancel something out...taking a step backwards

OpenStudy (anonymous):

i foiled the top but didn't add up yet, is this correct so far? 4x+9x+2x+6

OpenStudy (usukidoll):

\[(3x+2)(x+3) -2x^2 = 3x^2+9x+2x+6 = 3x^2+11x+6 -2x^2 = x^2+11x+6\]

OpenStudy (usukidoll):

remember \[(x)(x) \rightarrow (x^1)(x^1) \rightarrow x^{1+1} \rightarrow x^2\]

OpenStudy (usukidoll):

so... \[(3x)(x) \rightarrow 3x^{1+1} \rightarrow 3x^2\]

OpenStudy (anonymous):

Ohhh ok, so now i have x^2+11x-6/2x(x+3)

OpenStudy (anonymous):

would we cancel our the 6 from 3?

OpenStudy (usukidoll):

no

OpenStudy (usukidoll):

I don't think this can be factored... the only pairs for 6 is 6 x 1 1 x 6 2 x 3 3 x2 that's 6+1 =7 3+2 = 5 and subtracting will make it worse

OpenStudy (anonymous):

maybe thats the restrictions they talk about?

OpenStudy (anonymous):

i think

OpenStudy (usukidoll):

\[\frac{x^2+11x-6}{2x(x+3)}\] maybe I wish it was more specific.. there is also what values of x makes this undefined too

OpenStudy (usukidoll):

2x(x+3) = 0 split into 2 problems 2x=0 x+3=0 solving x ... X = 0,-3

OpenStudy (anonymous):

those are probably the restrictions Cx , i think it does mean undifined

OpenStudy (anonymous):

ok so the next question wants me to do this 

OpenStudy (usukidoll):

\[\frac{x^2+11x-6}{2x(x+3)} \times \frac{x+3}{x+1}\]

OpenStudy (anonymous):

its x-1 cx

OpenStudy (usukidoll):

we can cancel out the x+3... it's in the numerator and denominator \[\frac{x^2+11x-6}{2x(x+3)} \times \frac{x+3}{x-1}\]

OpenStudy (anonymous):

oh true! so they would disapear?

OpenStudy (usukidoll):

YUP! :D

OpenStudy (anonymous):

yes! the easiest step on here lol cx so then the problem would be x^2+11-6/2 times x-1

OpenStudy (anonymous):

*divided by 2x not 2

OpenStudy (anonymous):

how would that work? since the second one isnt a denominater nor a numerator?

OpenStudy (usukidoll):

\[\frac{x^2+11x-6}{2x} \times \frac{1}{x-1} \rightarrow \frac{x^2+11x-6}{2x(x-1)}\]

OpenStudy (usukidoll):

huh?! ok wait I just latex the fractions without x+3 because the numerator had one and the denominator had one .

OpenStudy (anonymous):

ohhh ok i remebr learning that step now, so it would stay like that? :o

OpenStudy (anonymous):

what happened :o

OpenStudy (usukidoll):

what is the next question?

OpenStudy (anonymous):

you said they cancled eachother out cx, and now i have to divide that equation by 6/x-1, but it also says to (Discuss why the degree of the resulting denominator did not change from your expression’s degree)

OpenStudy (usukidoll):

\[\frac{x^2+11x-6}{2x(x-1)} \] divided by \[\frac{6}{x-1}\]

OpenStudy (anonymous):

yea

OpenStudy (usukidoll):

if you flip the second fraction... surprise x-1 gets canceled out too.

OpenStudy (usukidoll):

\[\frac{x^2+11x-6}{2x(x-1)} \times \frac{x-1}{6}\]

OpenStudy (anonymous):

ohh and it turns into a mutplication by flipping?

OpenStudy (usukidoll):

when dividing fractions... you flip the second fraction... in the end you're multiplying

OpenStudy (anonymous):

Oh ok, that makes it easier cx

OpenStudy (usukidoll):

\[\frac{x^2+11x-6}{2x} \times \frac{1}{6} \rightarrow \frac{x^2+11x-6}{12x}\]

OpenStudy (usukidoll):

I think it's safe now to combine the denominator...

OpenStudy (anonymous):

oh yeah i keep forget that the single number would be on bottom and the 1 on top, and yea it does xD

OpenStudy (anonymous):

so then i would have to explain *Discuss why the degree of the resulting denominator did not change from your expression’s degree*

OpenStudy (anonymous):

what the flip does it mean by that lol :x

OpenStudy (usukidoll):

what is that?!

OpenStudy (usukidoll):

12x that's first degree isn't it?! awww c*** I haven't dealt with these questions in a while. pure math doesn't have this

OpenStudy (anonymous):

well we got the retrictions Q , and i just wanna finish this one assignemnt thingy lol, you look like your doing it right! cx its just these wierd Q's

OpenStudy (usukidoll):

how did it not change from the expression's degree? could it be referring to the original question?!

OpenStudy (usukidoll):

the first question you had ... contained 2x our fourth question had the denominator 2x(6) -> 12x I think they are the same first degree. That's just a guess ..